TEX and TUG NEWS

TEX and TUG NEWS is a newsletter for TEX and L^{A}TEX users alike: a forum for exchanging information, tips and suggestions; a regular means of communicating news items to one another; a place where information about TEX and TUG can be quickly disseminated. Throughout the newsletter “TEX” is understood to mean TEX, L^{A}TEX, AMS-TEX, and other related programs and macros.

This is a prototype issue of the newsletter. If the TUG Board of Directors decides to continue this publication, all the following will apply.

TEX and TUG NEWS is produced with the standard L^{A}TEX distribution, and is to be as portable a document as possible. This is done on purpose, in order to look at eventual electronic distribution of the source file. To test this, the source file of this issue will be placed in the aston, clarkson, heidelberg, and labrea archives initially.

The entire contents of this newsletter are being placed in the public domain. Copying and reprinting are encouraged; however, an acknowledgement specifiying TEX and TUG NEWS as the source would be appreciated.

As a newsletter for and by TEX users, extensive participation and comment will be solicited. Submissions to TEX and TUG NEWS will be encouraged. They should be short, the macros must work, and the files must run without special font or graphics requirements: this is to be a portable newsletter (the new font naming scheme has not yet been implemented). Correspondence may be sent directly via e-mail to: tug@math.ams.com, with the subject line NEWSLETTER. Regular mail should be addressed to the Publications Committee, c/o TUG Office at the address below.

Kind acknowledgement is made to Lance Carnes of Personal TEX Inc. and Peter Abbott of Aston University for providing complimentary camera copies of this issue of TEX and TUG NEWS.

TEX Users Group P.O. Box 9506 Providence, RI 02940-9506 USA Phone: 401-751-7760 FAX: 401-751-1071

                      E-mail: tug@math.ams.com

TEX is a trademark of the American Mathematical Society

Editorial

Welcome to TEX and TUG NEWS!

TEX and TUG NEWS is a TEX Users Group newsletter to supplement and extend TUG’s services to its members. After 12 years of TUG, it’s time to expand and change, a recurring theme in the past year, as evidenced in the board’s own composition, in the changes introduced into the annual meeting, and in the production of this prototype newsletter.

TEX and TUG NEWS is for TEX users to communicate with one another, and for the TEX Users Group to communicate with its members. It is intended to be a pragmatic source of information: an exchange of news items, columns and comments, letters to the editor, hints and tricks and ideas, the latest comings and goings in the community, books and articles of interest, and so on. As a “proto-Mission Statement,” this list is of course provisional.

The TEX and TUG NEWS prototype has two purposes: (a) to see what is involved in producing a newsletter; and (b) to get some important board communications/business out to the membership. This issue is atypical as it has a fair amount of category (b) material; these pages will be replaced by regular newsletter columns and articles starting with the first regular issue.

The TUG board has recognized the need for a newsletter for years and, if this prototype issue shows that a newsletter can be fiscally feasible, the board will approve the continuation of the project.

If that approval is given, lots of help will be needed. This is intended to be a users’ newsletter. You will be needed to contribute as much as you can, as often as you can, to get this newsletter onto a regular footing. Constructive criticism is always necessary and is invited; find fault and offer an improvement or an alternative. Style and format can only improve over time.

This publication is a prototype in format and in content — even its name may change. Potential items for future regular issues of the newsletter (a bi-monthly publication of 32 pages if we can get sufficient material from the thousands of TEX users out there …) might include: news items of course, regular and/or rotating columns, information exchange, help and advice, etc.

Submissions which have already found a place in this issue include an outline of a handy little L^{A}TEX add-on macro file, jeep, pulled together by Peter Stewart. There’s a teaser news item from Theo Jurriens on TEX in the Soviet Union — with more to come at the EuroTEX meeting in September; and an announcement on a new Greek TEX discussion group started up by Yannis Haralambous (Yannis will also be a discussion leader on this topic at the July 1991 TUG Meeting in Dedham).

There is a general report on the overall changes which have been effected in the organization (pp. 7–10), and summaries of board actions taken at the last two special board meetings (pp. 11–19). These provide a context for the letters which follow, published at the authors’ request.

The most important change affecting the membership is the change from an appointed Board of Directors to an elected board of 15 starting this year. The elections will be open to the entire membership, not just to those attending the Business Meeting at the Annual Meeting. The election procedures and a nomination form are to be found on pp. 25–26 (note that this supersedes information which appeared in TUGboat 12, no. 1).

Regular column editors are being sought, and I’ve written up a few potential columns as suggestions — one on TeXhax and another with some little tricks using \llap.

And there’s an Update on the 12th Annual TUG Meeting: be sure to read it (pp. 28–29), and make plans to join us in July. Other meetings and conferences which might also be of interest to the TEX-using community are listed in the Upcoming Events section.

So, please read on! Make notes in the margins. Volunteers, comments, suggestions, submissions will all be needed if we are to have a successful newsletter. Send all correspondence regarding TEX and TUG NEWS to tug@math.ams.com, with subject line NEWSLETTER. Participate! Bring comments and marked up copies to the July meeting. We’ll try to set up some kind of forum for talking about the newsletter — I don’t know how, as the schedule is very tight, but we’ll figure something out.

And if you haven’t renewed your TUG membership, or know someone who hasn’t joined or renewed: the back cover gives you eight great reasons to do so! Circulate the information — and help get more members!

This issue contains contributions from: Barbara Beeton, Regina Girouard, Yannis Haralambous, Doug Henderson, Theo Jurriens, Peter Stewart, and myself. Jackie Damrau and Will Ward receive my thanks for having provided the initial code for the \Section command. And I would like to express heartfelt thanks to Ron Whitney and the TUG office staff for providing material, advice, and reminders on various production aspects I was about to overlook. The prototype is the result of a truly collective effort.

Christina Thiele

Editor, Prototype TEX and TUG NEWS May 1991

jeep: A General Purpose Style File

Introduction

Jeep is a L^{A}TEX style file written to supplement the standard article and report styles. The name jeep is a continuation of Leslie Lamport’s analogy, in which he compares TEX to a racing car and L^{A}TEX to a family sedan. The modifications here convert the sedan to a general purpose vehicle — a jeep. The modifications, like the original jeep, are not pretty, but the casual L^{A}TEXer can cover a lot of territory with them.

It is important to keep in mind that jeep was not planned; rather it evolved as I expanded my use of L^{A}TEX. It has not changed much recently, and it has served me in a variety of applications. But being tailored to my needs, jeep will not necessarily meet all the needs of others.

Nor is jeep an original creation. I am no TEXpert, and my guiding principle has been: never create when you can adapt, and never adapt when you can copy. Thus, jeep is indebted to a number of people who took the trouble to write useful macros and distribute them to the community. Needless to say, the greatest debt is owed to Leslie Lamport for creating L^{A}TEX in the first place.

Features

Full descriptions and explanations on what jeep can do are contained in the file jeep.tex (see end of this article for details). A brief listing is provided here, so that readers can decide whether jeep is for them.

1. The default page size is changed to take full advantage of the default North American paper size of 8\frac{1}{2} × 11 in. The dimensions are satisfactory with eleven- and twelve-point type, not so pleasing with ten-point type.

2. Commands are provided to allow the user to change the formats of chapter, section, and theorem heads. The default section and subsection heads are made somewhat smaller than in the standard styles, and section numbers are followed by a period.

3. At the user’s option, equations and theorems may be numbered within sections or subsections.

4. The default page style is Lance Berc’s three-part header style. It provides for left, right, and center headers and footers, which may be changed at will by the user. The style has been extended to make it useful in creating headers for books.

5. There is a command to alter the page style to put a rule under the headers. The rule may be suppressed on individual pages (usually the first).

6. The default numbering of the enumerate environment has been changed to Arabic at all levels.

7. The outline environment, a variation of the enumerate environment, puts less space between its items. It is especially suitable for long, nested lists of short items, such as a course outline.

8. The \tighttoc command produces a compact table of contents, useful for long journal articles.

9. A \symbolnote command produces footnotes marked by the conventional symbols rather than numbers (see L^{A}TEX Manual, §C.7.4). These footnotes are independent of the usual sequence of numbered footnotes, and are especially useful in creating title pages.

10. A symboleqn environment creates a displayed equation with “equation number” specified by the user. These equations are independent of the usual sequence of numbered equations. They are useful for referencing local equations in problem sets.

11. Modifications by John Hobby make the TEX commands \bigl, \bigr, etc. compatible with L^{A}TEX.

12. A literatim environment, a variation of Lamport’s alltt environment, allows commands in a verbatim-like environment.

Availability

jeep consists of two files: jeep.tex (the documentation) and jeep.sty (the style file). They are available by anonymous ftp from:

• sun.soe.clarkson.edu in the directory pub/tex/latex-style
• thales.cs.umd.edu in the directory pub/jeep

P.W. Stewart

University of Maryland

News from Around …

TEX in the Soviet Union Theo Jurriens

Rijksuniversiteit Groningen

Since 1983 I have visited the USSR eleven times. Such visits always give rise to a boom in contacts. One long-standing contact is Dr. Anatoli Urvantsev, former Dean of the Faculty of Mathematics and Computer Science of Novosibirsk University, currently working at the Computing Center; the University is twinned with Stanford University. In November, he spent two weeks in the Netherlands and also attended an NTG meeting in Utrecht. At that meeting he opened his TEX horizons. Thanks to e-mail our contacts have increased and I volunteered to go to Siberia to give a course.

Urvantsev selected 13 students (11 female, 2 male) from the University Computing Center and from a scientific publisher, ranging in age from 20 to 50 (!). Although he promised a good level of English, the reality was a bit different. But thanks to my Dutch friend and translator I survived.

In 6 three-hour sessions I presented the basics of L^{A}TEX, based on the principle of teaching by example. In total, my students worked on 40 exercises. And to my own surprise we succeeded in doing everything.

In groups of two or three students we used ATs to do our job. ArborText was kind enough to supply a full version of μTeX. This version stayed behind and I also left behind a Unix tape with the latest versions of TEX and . This tape is now touring the Soviet Union.

So, to stimulate TEX it’s necessary to first have an organisation, like CyrTug. In the Soviet Union having enthusiastic people is not enough — you have to organise. Just before the deadline for this newsletter I heard the good news that CyrTug will be meeting May 22–25, in Moscow.

After my return I helped my Siberian friends solve their TEX problems via e-mail. By the end of the year I hope to return, to give seminars in Leningrad, Krasnojarsk and Yakutsk. But before then, I will be giving a more detailed report on my Siberian trip at EuroTEX in Paris (Sept. 23–25, 1991).

New LISTSERV for Greek TEX Users Yannis Haralambous

Université de Lille

A new e-mail discussion list ELLHNIKA has been opened at DHDURZ1 on Bitnet, thanks to DANTE e.V. and especially Joachim Lammarsch. This list is intended to cover the following topics:

1. Until the fall of 1991, discussion on a standardization of Greek TEX. The two main areas of concern are transliteration of Greek to the Latin alphabet and establishment of a standard Greek font table for TEX.

2. General discussion on the relationship between TEX, Modern Greek Language and/or Greece.

3. General discussion on the use of TEX in the world of scholars interested in Ancient Greek.

The languages used in the list are English and (Modern) Greek. A biweekly information bulletin will among other things make an inventory of proposed transliterations which can then be used for messages sent in Greek.

The final goal is to make a complete Greek TEX, both for Ancient and Modern Greek with the corresponding L^{A}TEX styles, hyphenation patterns and font families. Since discussions on the standardization are open from now until the fall of 1991, every TEX user is encouraged to state his/her ideas and opinions by joining our list!

Σα̃ς περιμένoυμε!

   To subscribe send SUB ELLHNIKA to LISTSERV@DHDURZ1.BITNET.

The Truth about TEX Christina Thiele

Carleton University

Have you ever run across articles or descriptions of TEX and/or its related programs, and thought the text wasn’t quite right? It happens. And TUG would like to do something about it, by providing the author or publisher with more accurate information. So if you do find such items, please send full details to the TUG office, c/o the Publications Committee, or to tug@math.ams.com, subject line Publications Committee. Provide the correct title of the item and where it was found (a photocopy of the item would be ideal), along with a brief description of the context in which it appeared (an article on word processing, comparison of page makeup programs on PCs, etc.).

A Message from the Board of Directors

The intent of this note is to help bring into focus the board’s actions to provide new leadership and direction for the TEX Users Group.^{1}

In the past, very little has been reported to the membership about the officers and board of directors, what they do, and how they serve the Users Group. This letter is intended as a start to changing this practice, and relates some specific changes that have occurred recently.

At present the board of directors is composed of the following individuals:

Nelson Beebe David Fuchs  Doug Henderson Joachim Lammarsch
Barbara Beeton Richard Furuta Alan Hoenig  Pierre MacKay
Lance Carnes Bernard Gaulle Patrick Ion  David Ness
Bart Childs Regina Girouard Cal Jackson Craig Platt
Malcolm Clark Roswitha Graham David Kellerman Christina Thiele
John Crawford Dean Guenther Donald Knuth Kees v.d. Laan
Allen Dyer  Hope Hamilton David Kratzer Hermann Zapf

The TEX Users Group articles of incorporation and the bylaws are the foundation of the organization; they form the essence of the Users Group, in both a legal and organizational sense. They include the descriptions of the various groups that make up the Board of Directors, including the Executive and Finance Committees, and some detail as to the specific charges to committees. Since their adoption, significant changes have been made to the bylaws so that they reflect actual practices of TUG or address its future goals.

New focus

As a piece of software, TEX is one of the most ambitious and exciting large-scale projects for typesetting known. It is used actively worldwide. It offers users unprecedented sophistication and quality in typesetting. In its initial form it was outstanding, and in 1989, its creator, Donald E. Knuth, decided to enhance it, to allow it to be used as a multilingual vehicle, extending the range of the input character set to double the size. This work is largely due to the vision of one man, Knuth, who saw fit to improve on the already excellent typesetting software. This vision of TEX being used throughout the world was already in place before the changes were made, and the changes ensure that the software environment known as TEX will remain solid in its abilities to give consistent and undeniably excellent typesetting well into the future.

1This report was written by Doug Henderson, in consultation with other members of the Board of Directors.

It is interesting to note the parallel between taking something that was good and making it better with respect to our Users Group. The board of directors feels that with the growth in the use and uses of TEX it is time to change from a user group that is good to one that is better. To accomplish this we need a clear vision of what it is that the Users Group is to do. The board has one disadvantage when trying to define the role of TUG, because the size of the board means there is a lack of clarity that results from a single viewpoint. Instead there are roughly a dozen or more viewpoints (not all of the 26 board members are actively participating) struggling to find worthy, obtainable goals. However, this disadvantage can also be considered an advantage because with a large board, representing many different types of users and many different types of systems, we can be sure we are considering goals that should benefit the entire TEX community.

In recent years the TEX Users Group has not been adequately prepared to respond to new user needs resulting from continually expanding uses of TEX throughout the world. The level of participation from our officers and board of directors had declined, and as a result, some stagnation had occurred. To help us focus, as a group, on relevant issues both now and in the future we need to embody our ideas in the form of a Mission Statement. This will allow us to focus on issues that will be supportive of such an overall set of goals, with a better understanding of how to accomplish them.

The summaries of the Cork meeting in 1990 and the special Dedham board meeting in March of 1991 show that the board’s work is not easy (see next article). They also show some of the good ideas for making our Users Group better. Some of the ideas that have evolved include selecting an individual who can work as an evangelist for TEX throughout the world as our next executive director. We will begin the search for this person as soon as possible. It will be an open selection process inviting all who feel qualified to step forward and claim that their goals of leadership are the same as the needs of the organization. Spreading the word about TEX to others and working with the world community to further expand the use of TEX are just two such goals.

Some changes

One positive step taken by the board was the formation of a Planning Committee. It is anticipated that this group will come up with a number of suggestions to help our organization keep pace with the promotion of TEX and the TEX Users Group while remaining true to the ideas put forth in our mission statement.

Another significant advance for our Users Group was the restructuring of the election procedure to a more democratic process. In the near future, the board of directors will be smaller and it will be an elected body. The voting will take place by a standard mailing to the entire membership. The board felt it was time to move away from voting for officials only at the annual meetings (where at most 5 to 10 percent of the members could participate), to a more open format where all TUG members will be encouraged to vote. In light of this new revitalization of elections, members should keep in mind the significant amount of work that will be required to serve on the board in the future.

Changes have occurred over the last five years that have moved TEX out of the exclusive domain of academia to being accessible to any person with a personal computer. As such, we need to explore ways to reach a very large group of people (roughly 10,000 to 15,000 personal computer users) to encourage them to join our Users Group, which currently totals about 4,000. One way to do this is by offering them advantages that are of interest to TEX users, rather than just to implementors or computer programmers. The quote from Gerald Oster in Optical Art 1965 is relevant. “Computer technology has far to go to duplicate the spirit and excitement of art. When such a day is reached, however, it will require an artist to do the programming.” This quote was reproduced in TUGboat in 1987 (Vol. 8, no. 1), but is pertinent today for our Users Group. We have the stage set for artists, publishers, and a host of others to start using this incredible typesetting technology. We need to find a way to bring it to them in digestible, less technical terms.

A second information forum, in the form of this newsletter, aims to help bridge the gap between the highly technical and the pragmatic needs of TEX users. TUGboat is something we can all be proud of as a scholarly publication for TEX programming and development. Its excellence is due in no small part to the vision and energy of Barbara Beeton and a host of authors and editors. But making TEX usable to a larger segment of the world is also a worthy goal.

We also need to encourage macro development projects such as Frank Mittelbach’s reworking of L^{A}TEX, which removes much of the difficulty of using TEXin-the-rough, so to speak. It is also clear that software developers on personal computer platforms can have a lot to do with the acceptance of TEX by making TEX faster, easier to use, and less complicated to install, and by integrating the different elements of the typical TEX processing cycle—encode, TEX, preview, print—in one package.

It is also time to find the energy to promote the public domain versions of TEX as well as the commercial flavors. With this in mind, the board has recently appointed Jon Radel as our Public Domain Software Coordinator. Jon has been working in an unofficial capacity for some time now, and we appreciate the work he has done and will continue to do for us in the future. Recently, Jon was given a disk duplication machine to help him with his public domain distribution.

Originally, our Users Group was a place for developers to present and exchange ideas about implementing TEX on different machines. Without the past foundation of solid research and development, we would not be at the point we are today, poised for carrying the work further to present TEX to the rest of the world — those who are not computer programmers. We are, after all, the TEX Users Group. This is not to say that development work should stop, or be given a diminished role in our Users Group. Instead, we should explore good uses of all the various talents within our group for their particular skills. For instance, one thing that would ensure the continued growth of TEX in the near to distant future is to encourage our academic colleagues to become involved with standards committees. We have identified the need for TUG to be involved with font-naming conventions, understanding of and synergism with markup languages such as SGML, and interaction with ODA (Open Document Architecture) and ODIF (Open Document Interchange Format) standards committees. This activity would ensure that TEX is fairly represented in these areas and does not become obsolete by isolation from evolving standards. We should take this to heart and get involved now.

Although there are a few dissenting voices among the board of directors, as evidenced by the resignation of three members who no longer wished to continue on the board, the spirit for the future of TEX and TUG is very good now. With the changes already mentioned, as well as ideas from you, the membership, we can all enjoy the new structures put in place recently, and begin spreading the word. It’s also time we started interacting more with other user groups such as UK TEX Users Group, Dante, and Gutenberg in a cooperative spirit.

The word is TEX, and the various world Users Groups are the vehicles.

Summary of TUG Board of Directors Meetings

In recognition of the amount of work that must be done as TUG enters its second decade, the TUG Board of Directors held two special sessions after its regular meeting at the June 1990 meeting in Texas. The first special board meeting was held in Cork, Ireland, in September in conjunction with TEX90; the second meeting was held in March 1991 in Dedham, Massachusetts. The meetings reflected the feeling that as TEX and TUG have grown, the needs of the Users Group are changing and the role that TUG plays must also change. Many of the motions passed at these meetings show the Board’s responses to those changes. What follows is a summary of the actions taken at those two meetings.^{2}

Board of Directors Meeting, Sept. 1990

This meeting was held on September 8 and 12, 1990, at University College, Cork, Ireland. Directors in attendance were Nelson Beebe, Barbara Beeton, Malcolm Clark, Allen Dyer, Bernard Gaulle, Regina Girouard, Raymond Goucher, Roswitha Graham, Dean Guenther, Doug Henderson, Alan Hoenig, David Kellerman, Joachim Lammarsch, Kees van der Laan and Christina Thiele. The directors who could not attend this meeting were: Lance Carnes, Bart Childs, John Crawford, David Fuchs, Rick Furuta, Hope Hamilton, Patrick Ion, Cal Jackson, David Kratzer, Pierre MacKay, David Ness and Craig Platt.

TUG’s Organizational Structure

At the Cork meeting the board voted on several changes to TUG’s operating structure to allow for full board participation in TUG’s affairs and to provide members with information about board and TUG activities.

First, it was agreed that it would not be appropriate to have a TUG staff member be a member of the Executive committee (EC) and Board of Directors. These bodies could certainly not function effectively without the presence of the Executive Director (ED), but it was agreed that the ED should not be a voting member of either body. Two motions were made by Allen Dyer. The first would remove the ED from membership on the EC (and by default from the Board of Directors) and the second would amend the bylaws to make it clear that the EC works in the place of the board and that it should be responsible to and report to the board. The first was seconded by Christina Thiele and passed on a 9 – 4 vote (9 votes in favor, 4 votes against). The second motion was seconded by Regina Girouard and passed on a 13 – 0 vote.

2The following summaries were written by Regina Girouard, in consultation with other members of the Board of Directors.

According to the TUG bylaws, members of the Executive Committee (EC) should be members of the board. However, because the members of the EC are all officers, they are ex officio board members. The same is not true for the Finance Committee (FC), however. Two members of the FC are appointed by the board and are, therefore, not ex officio members of the board. To rectify this situation Allen Dyer made a motion to make all members of the Executive and Finance committees members of the board. The motion was seconded by Christina Thiele and passed unanimously (with two members abstaining).

A great deal of time was spent discussing the structure of the board of directors and how the board can best operate effectively. One of the problems the board addressed was how to allow the board ample time to discuss and act on the many matters it should be involved in. With TUG having an international board and with there normally being only one board meeting per year, it is difficult to accomplish all the work that is necessary. To make board work easier Allen Dyer made a motion that would allow the board to conduct its business by mail, either e-mail or regular mail. Motions can be proposed and voted on by mail. Mail business would require a two-thirds vote of the board for approval rather than the usual simple majority. This process would normally be used for routine matters and will allow the in-person board meetings to concentrate on the more important items. The motion was seconded by Barbara Beeton and passed unanimously. It was agreed to postpone the discussion about the size of the board until the next board meeting.

The TUG board often establishes ad hoc committees, in addition to those which are permanent and outlined in the bylaws. To provide some guidance and structure to the ad hoc committees, Christina Thiele proposed a motion on basic procedures: the committees should report the names of their members, their chair, provide a statement of purpose, and establish a reporting schedule. The motion was seconded by Doug Henderson, and passed by a 12 – 0 vote (two members abstaining).

The board was concerned that the work of the board and staff not appear to be secret. To provide an open atmosphere, Allen Dyer proposed that the records of the TUG office (except for those protected by federal or state law) be open to all TUG members. The motion was seconded by Doug Henderson and passed on an 8 – 2 vote (3 members abstained.) As well, David Kellerman made a motion that TUG publish the minutes of its Board of Directors meetings so the membership will be aware of the work being done by TUG and the problems they encounter. His motion also called for the minutes to be made available electronically once they have been approved. The motion passed unanimously.

Allen Dyer proposed that TUG adopt a “Minutes and Procedures” book, in both paper and electronic form, to help provide for better management and training, and to provide an accurate record of TUG policies and decisions. The motion was seconded by Regina Girouard and passed unanimously.

There had been much discussion at the Texas meeting about the role TUG should be playing in the TEX community and whether the interaction between the various components of the current TUG structure (the staff, the officers, the board, the Executive and Finance Committees) was the best way to meet the growing needs of TUG and TEX. There was a great deal of concern among board members (and the membership) about the future of TEX. However, it was felt that before these concerns could be adequately addressed, the board must first address its ideas of the goals of TUG and how to be sure the various components of TUG are working towards those goals.

Many of the motions passed at Cork reflected the feeling of many of the board members that, perhaps because of the growth of TEX and TUG, the operating structure that had worked for TUG in the past was no longer the best method. Until this year the Board of Directors met only once a year. Annual meetings could not, however, provide the ED with the advice and consent that is essential to operating an organization. For this reason TUG had two smaller groups (the Executive and Finance Committees), which would meet with the ED on a more regular basis. According to the bylaws, the board is responsible for the operation of TUG. However, since the ED worked primarily with the Executive and Finance Committees, the current structure left the board unaware of many of the problems facing TUG and of some of the decisions made for TUG. Attempts had been made in the past to try to ensure that the board would be fully involved in these problems and decisions, but these attempts had not been successful.

After a very lengthy discussion it appeared that there was a basic disagreement about how TUG should operate and what direction it should be going in. The ED was satisfied with the way TUG had always operated, i.e., he worked primarily with and for the Executive and Finance Committees, and the board should not be concerned with all the business details of TUG. In fact, the board was not to be allowed access to some of the information about the operation of TUG even though they were legally responsible for all of the actions of TUG. In addition the board wanted TUG to start moving in new directions, becoming more involved in L^{A}TEX and TEX on personal computers, and taking more of a leadership role in the TEX community by means of standards or reviews of new TEX software. However, the ED believed in maintaining the status quo. After hours of discussion the board decided that the contract of the Executive Director would not be renewed when it expired at the end of 1990. The board charged the Executive Committee to develop an appropriate job description for the ED position, and to prepare a plan for the initiation of an open selection process. A motion to that effect was made by Allen Dyer, seconded by Regina Girouard, and passed on 9 – 3 vote (with one member abstaining).

The board was very appreciative of the work the ED had done for TUG. They acknowledged that TUG would not have been able to grow (in both size and services offered) without his efforts. In addition to a generous compensation package, the board voted to make him an Honorary Member of the board as the Founding Executive Director of the TEX Users Group.

The board regrets that this action was necessary but feels that it is essential to the future of TUG that the chief executive officer work in conjunction with the board to move TUG forward into the 90’s.

TUG as a Service Organization

Some of the other motions passed reflected TUG’s goal to be as much of a service as possible to the entire TEX community.

Barbara Beeton proposed that TUG remove the postal surcharge imposed on members outside of North America for their subscriptions to TUGboat. While it does actually cost more to mail copies outside of North America it was felt that charging non-North Americans more did not agree with TUG’s statement that it is an international organization trying to meet the needs of users throughout the world. The motion was seconded by Doug Henderson and passed unanimously (with one director abstaining).

Malcolm Clark moved that TUG’s Scholarship Committee be composed of two TEXperts and the previous scholarship winner. This motion was seconded by Dean Guenther and passed unanimously.

There was a lengthy discussion of the roles of User Groups, the different needs of the user groups in various countries and/or language-based user groups, and whether TUG could become an International User Group that could respond to all these varying needs. Dean Guenther proposed that the President establish an International TUG committee; the committee should report to the board with recommendations for identifying and remedying the problems in serving the international community. The motion was seconded by Bernard Gaulle and passed unanimously (with two directors abstaining).

There was agreement that TUG should be doing all it can to help spread the word about TEX. Nelson discussed his new TEX lists: texbookn. These are lists of publications about the use of TEX and publications produced with TEX. The board is considering a public relations campaign that could include producing a catalog of available TEX products and mailings to universities to describe the advantages of TEX. Allen Dyer made a motion that TUG purchase a disk duplicator to help provide a low cost distribution method for public domain software. This motion was seconded by Doug and passed 12 – 1. Allen and Doug also suggested that the board authorize up to $5,000 to advertise TEX and TUG in a microcomputer journal such as Byte. Since the return from such advertising was unclear, the board felt that, for the time being, there were more effective uses for the money, and the motion was defeated on a 4 – 8 – 1 vote. The board considered some objectives for TUG in 1991–92 but since there was not sufficient time to go into great detail, this subject will be considered via e-mail discussions and during the next board meeting.

There was a lengthy discussion about the need for certification of TEX-related software. One suggestion was that software developers be given a certification form stating that their products have passed the trip and/or trap tests. They would have to sign, notarize and publish this certification. If it turns out that their product does not pass these tests, purchasers would have the invalid certification to use against the developer.

The meeting ended two very long days of discussions and decisions regarding TUG’s operation, and its role in the larger TEX community.

Board of Directors Meeting, March 1991

This meeting was held on March 2–3, 1991 at the Hilton hotel in Dedham, Massachusetts. Directors in attendance were: Barbara Beeton, Lance Carnes, John Crawford, Allen Dyer, Bernard Gaulle, Regina Girouard, Doug Henderson, Hope Hamilton and Christina Thiele. Also present at this meeting were TUG staff members Ron Whitney and Cliff Alper. The directors who could not attend this meeting were: Nelson Beebe, Bart Childs, Malcolm Clark, David Fuchs, Rick Furuta, Roswitha Graham, Dean Guenther, Alan Hoenig, Patrick Ion, David Kellerman, David Kratzer, Kees van der Laan, Joachim Lammarasch, Pierre MacKay, David Ness and Craig Platt.

Organizational Matters

With nine board members present there was a quorum so this was an official board meeting. Normally the President presides at board meetings; however, since Nelson Beebe could not attend, it was agreed that Hope Hamilton would chair the meeting. As chair, Hope did not vote on any of the motions but she was an active participant in the discussions.

In recognition of the fact that many board members could not attend, one of the motions passed at this meeting was to delay for 30 days the implementation of any motion passed. The draft minutes of this meeting were to be distributed to the full board immediately after the meeting. The 30-day delay allowed time for all board members to review the actions taken at the March meeting and, with a simple majority vote by e-mail, nullify any of those actions. The 30-day period passed with no dissenting votes being cast so all the motions approved at this meeting are now valid.

Several motions were passed that dealt with the TUG office. First was a motion by Allen Dyer (seconded by Regina Girouard) to initiate work on a written set of “TUG Operating Procedures.” These Operating Procedures, which were authorized at the Cork meeting, will provide for better management of TUG by (1) supplying guidelines and formal statements of TUG policy; (2) providing for continuity of operations regardless of personnel changes; (3) providing for more rapid training of personnel or volunteers; (4) helping to maintain consistent decision making; and (5) giving the membership of TUG an opportunity to see how the affairs of the corporation are conducted. The motion passed unanimously.

An old idea that was on the agenda for the Texas board meeting but never discussed was revived by Regina Girouard. This was a motion that TUG waive the meeting registration fee for any board member attending a board meeting, which is being held in conjunction with a TUG conference, in a location outside their own continent. It is hoped that this waiver of a meeting registration fee might make it more financially feasible for more board members to attend distant meetings. The motion was seconded by Bernard Gaulle and passed unanimously.

Allen Dyer made a motion that up to $6,000 of the budgeted amount for committee expenses be authorised for reimbursement of expenses incurred for this board meeting. The motion was seconded by Bernard Gaulle and passed unanimously.

A subcommittee of board members met on Saturday night with Ron Whitney, TUG’s Acting Business Manager, to discuss some personnel matters that had been pending for some time. On Sunday the board heard the report of the subcommittee and, on a motion made by Regina Girouard and seconded by Doug Henderson, voted unanimously to approve a few overdue and promised salary increases and/or benefit changes for the TUG staff. The board also appointed a subcommittee of Hope Hamilton (chair), Lance Carnes and Allen Dyer, to negotiate a contract for a TUG Business Manager. This contract will be for approximately one year and will be approved by the full board at the July annual meeting. The board must have someone in authority pending the establishment of procedures for defining and selecting an Executive Director. This action will allow the board sufficient time to accomplish the task.

A great deal of time was spent discussing the structure of the board of directors and how the board can best operate effectively. E-mail voting, instituted at the Cork meeting, required refinements. Doug Henderson proposed that, for the board or any committee, e-mail motions would be formally made and seconded. Once seconded, there would be a two-week discussion period, followed by a two-week voting period, after which the votes would be tallied. The total number of votes cast (either Aye, Nay or Abstain) have to equal at least 50% of the total number of members eligible to vote, otherwise the motion fails. If the motion passes the 50% participation requirement it requires two-thirds of the votes cast in order to pass. The motion was seconded by Bernard Gaulle and passed unanimously.

There was one multi-part motion made by Allen Dyer (seconded by John Crawford), that dealt with TUG officers and the Board of Directors. the purpose of the motion was to make the TUG board more effective and more representative of the entire membership.

Some of the sections of the motion dealt with how people become members of the TUG Board of Directors. Currently, all members are on the board either as ex officio members or they have been appointed by the board. Some members have been appointed because of their role (site coordinators, TUG officers), and others were appointed because of their knowledge or TEXpertise. The board felt that this situation results in a board that is too large to function efficiently. In addition, since there is no term of membership, it is not clear whether an ex officio board member should (or does) remain a member of the board when he or she no longer holds that office.

To solve these problems it was decided to amend the bylaws so that future boards will consist of the TUG President, some number of elected board members, honorary members and non-elected Vice Presidents (usually heads of large language-based User Groups). The honorary members will be advisory; they will not be voting members of the board.

The election process for board members will be by a written mail ballot of the entire membership. TUG members may have theirs name placed in nomination by submitting a petition signed by two (2) other members, or any member may be nominated at the annual TUG business meeting [see elsewhere in this newsletter for information on procedures and the nominations form –Ed.].

Another section of the Dyer/Crawford motion dealt with the officers of TUG. The motion amends the bylaws to provide that the President of TUG will be elected by written mail ballot of the entire membership. The nomination process will be the same as that described above for election to the board. The other officers of TUG (Vice-President, Secretary and Treasurer) will be members of the board and will be appointed by the board. The board may still continue to appoint special Vice-Presidents or assistant officers if it deems that necessary.

The final TUG structural questions addressed by the Dyer/Crawford motion dealt with the TUG committees (Executive and Finance), that had worked as a ‘small board’. The bylaws were amended so that the Finance Committee has been dissolved and the Executive Committee will consist of the President and three other board members. The Executive Committee will be responsible for adopting interim procedures and policies (subject to approval by the full board), on behalf of the full board.

The motion also established a Technical Council which will consist of Site Coordinators, Wizards and other active TEX contributors. After the initial creation of the Technical Council by the board, the Council itself shall determine its composition and operating procedures, and its own purposes and goals; however, these should be consistent with the purposes and goals of TUG. The Technical Council shall designate a representative to attend meetings of the TUG board in an advisory capacity.

The implementation of these motions will not take place immediately. The next President of TUG will be appointed by the board at the 1991 annual meeting; after that future Presidents will be elected by the membership. The election of board members will take place after the 1991 annual meeting. The current Board of Directors will continue until that election takes place.

After long discussions, the entire Dyer/Crawford motion passed unanimously. An Elections Procedures Committee was established, to develop procedures

for TUG elections for Board Members and the TUG President, and to develop a plan for implementing elections in 1991. Those who have accepted an appointment to this committee are: Barbara Beeton, Dean Guenther, Alan Hoenig, David Kratzer, Pierre MacKay and Craig Platt. The motion was proposed by Allen Dyer, seconded by Christina Thiele, and was passed unanimously.

Allen Dyer also made a motion (seconded by Regina Girouard) which sets the number of elected board members at 15. That motion also passed unanimously.

TUG and the Future of TEX

Some motions were passed that dealt with the role of TUG in the future of TEX. First was a motion by Doug Henderson (seconded by Allen Dyer), that TUG should have a written Mission Statement so everyone in TUG will know exactly where TUG is going and what its role in the TEX world is. The motion charged a committee composed of John Crawford, Hope Hamilton and Christina Thiele, to seek input from the full board and to have a draft of a TUG Mission Statement ready for approval at the board’s meeting in July. The motion passed unanimously.

Next the board considered the formation of some new TUG committees. However, they first reviewed how TUG committees are formed, how they are charged, and how they operate. They agreed that in the future possible committee members will first be advised of their nomination to the committee along with what the committee is charged to do. A nominee’s appointment to a committee will not be official until the committee member has agreed to the appointment. Each committee will be charged with some form of regular reporting with the minimum reporting procedure to be a report presented at the annual meeting. Committee members will serve until (1) they resign from the committee; (2) they are removed from the committee by a vote of the Board of Directors; or (3) until the 1991 TUG meeting. In the future, to ensure active committees, the membership of all TUG committees will be reviewed at each annual meeting.

After these committee procedural matters were settled, the board discussed two new committees. On a motion made by Allen Dyer and seconded by Regina Girouard, the board created a TUG Publications Committee. The charge to the committee is: (1) to determine the goals of TUG with respect to publications;

(2) to review the TUGboat goals, policies and content to determine whether it satisfies the audience and what improvements might be made; (3) to examine the records of publications now offered for sale by TUG to determine what members and customers find useful; (4) to review available records to see what kinds of publications have been requested; (5) to compile a list of topics that are appropriate and useful and to undertake a search for authors and solicit manuscripts on those topics; (6) to develop a pricing policy consistent with agreed-upon goals that will make publications both affordable and self-supporting; (7) to develop a policy on what kinds of publications should be made available electronically; and (8) to develop a plan for establishing a newsletter. It was agreed that the newsletter should have the following characteristics: (1) it should be the root newsletter of the worldwide TEX community; (2) it should contain news, minutes, announcements and the like; and (3) it should adhere to portability so that it can be modified by a LUG into a local edition with respect to either contents or language. The motion passed on a 7 – 1 vote.

Those who have accepted an appointment to the Publications Committee are: Barbara Beeton, Malcolm Clark, Allen Dyer, Regina Girouard, Kees van der Laan, and Ron Whitney.

Christina Thiele made a motion to set up a temporary team of board members to produce a prototype issue of a TUG newsletter. Christina will be the editor of the prototype and will have it ready for distribution before the annual meeting. The board authorized the expenditure of up to $5,000 for printing and mailing the prototype. The motion was seconded by Barbara Beeton and passed unanimously.

Next, on a motion made by Allen Dyer and seconded by Barbara Beeton, the board created a TUG Planning Committee. The charge to the committee is to develop recommendations for TUG strategic goals and to develop a plan for implementing those goals. Those who have accepted an appointment to the Planning Committee are: Malcolm Clark, David Kellerman, Kees van der Laan, Pierre MacKay and Sam Whidden. The motion passed on a 7 – 1 vote.

A motion by Allen Dyer (seconded by Lance Carnes and passed unanimously) established a Scholarship Committee to continue the Donald E. Knuth scholarship award, and budgeted $2,000 for the award. Those who have accepted an appointment to the Scholarship Committee are Barbara Beeton, and Kees van der Laan.

Allen Dyer made a motion that TUG appoint Jon Radel the TUG PC Public Domain Software Coordinator. The PC Public Domain Software Coordinator is charged with accumulating and organizing TEX-related software in a format that can be widely distributed among PC users. He will also provide master copies of the accumulated software to the TUG Office on a regular basis. TUG will partially reimburse the Coordinator for expenses incurred in this work. This motion, along with the board’s decision to purchase a disk duplicator for the TUG office, will enable the TUG office to distribute Public Domain software to TEX users at minimum cost.

The meeting adjourned, having attending to all but a few items on the agenda, and set the next board meeting to begin July 13th, 1991, before the 12th Annual Meeting of the TEX Users Group.

20  Vol. 0, No. 0, May 1991  TEX and TUG NEWS

Communications

Re: Resignation as TUG Treasurer 26 December, 1990 To Nelson Beebe Dear Nelson

I resign as Treasurer of TUG, effective as soon as you can appoint a replacement. I have sent copies of this letter to Barbara Beeton, with a request to publish it in a forthcoming TUGboat, and Ron Whitney, for deposit in the TUG ‘Archives’ in Providence. I enclose a letter to the membership setting forth the reasons for my resignation.

Regards David Ness 803 Mill Creek Road Gladwyne PA 19035

Dear TUG Members Re: Resignation as Treasurer of TUG

The actions taken in the name of the Board of Directors of the TEX Users Group since June [1990] have, in my opinion, been both irresponsible and unconscionable. My deep disagreement with the actions taken by a dissonant plurality of this group necessitates my resignation from the Board of Directors and as Treasurer of TUG. I have resigned in a letter sent today to both the Executive and Finance Committees of TUG and to the TUG membership via the office in Providence. I have requested that this letter be published in TUGboat, along with a copy of my official resignation. Some Background First, I should note that other than the officers, members of the Board have never been elected to their jobs. Indeed, many of the dissonant plurality are relatively recent appointments to the Board. While this does not disqualify these few members from concern for the organization and its well being, neither does it particularly qualify them to speak on behalf of the organization as a whole. However, in my reading of the current bylaws of TUG there is little question that they can ‘legally’ proceed in the direction that they currently appear to wish to pursue. As a duly elected officer, my only recourse is to decline to join them on this journey by resigning my office.

Second, if there were problems with the administration of TUG, then in all good conscience they were problems of the Executive and Finance Committees. To the best of my knowledge and belief, at no time did the TUG Office in Providence act in any manner contrary to that which had been specified—nay, encouraged—by those of us on the Executive and Finance Committees. If the membership did not like the way that TUG was managed, and expressed that fact to the Executive and Finance Committees, then it was truly our fault that we did not communicate this to the Providence staff. If the membership did not like the way that TUG was managed, and did not express that fact to us, then blaming the Providence staff is unfair. I am proud, pleased, and honored by my association with the Executive and Finance Committees, so I concede no fault— intentional or unintentional—here on our behalf.

Third, the dissonant members of the Board are responsible for a number of actions. It is my belief that their actions have already cost TUG substantially more than $50,000. Indeed, I think that $100,000 would be a closer estimate. This is money TUG can ill afford and is, I believe, a poor choice of expenditure on the part of an organization which I expect may be under increasing financial pressure in the immediate future. As Treasurer, I refuse to collaborate with this egregious disregard for

TEX and TUG NEWS  Vol. 0, No. 0, May 1991  21

the financial well-being of our organization. However, since I believe that that Board does have the legal power to make these mistakes, I have no recourse other than to resign.

Fourth, I believe that the actions by the Board have succeeded in making TUG ungovernable. I do not think it likely that a useful consensus of more than 25 people can effect day-to-day management of an organization, particularly via electronic mail. It does not seem sensible to me to expect an office staff, including an Executive Director, to be expected to respond to demands made by a group of this kind.

In addition to these unfortunate actions, the communications by some of the members of this dissonant group have been in extremely poor taste and certainly not in the spirit of the organization that I was pleased to join many years ago. For example, I find the arid remarks written in ‘answer’ to a letter by some past officers— all of whom spent many years working for this organization—to be in bizarrely poor taste.

Fifth, I would suggest that the other members of the Board of Directors and the membership at-large be very wary of this small dissonant Cabal. I remember the last time I successfully engineered a coup d’état (it was in junior high school). We were quite successful—as this Cabal has apparently been—in taking over an organization. However, the coup d’état proved to be a coup de grâce, and the organization in question did not survive the year of our takeover. We won a battle, not a war.

I hope that the non-dissonant members of the Board, including those members of the Executive and Finance Committees who have the courage and the interest to see beyond this battle will reassert their strength and intelligence—and thus win this war. Otherwise, I suspect that this Cabal will insure a similar fate for TUG to the one I saw so long ago. Indeed, this may already have been anticipated by Lance Carnes who, as a leader of the Cabal, said of TUG (after the June Board meeting in College Station, and perhaps in the heat of winning a battle there) “Let’s change it or kill it!” I thought it a presumptuous remark then, but I now fear that he and his associates have managed to accomplish both of these objectives. No credit to them. And no help to TUG.

Some Thanks I have neither the energy nor the interest to fight this battle. It has drained my volunteer spirit. I withdraw from TUG with profound thanks to many people. First, I owe gratitude to my colleagues and former colleagues on the Executive and Finance Committees. In my experience with boards of commercial organizations and academic institutions, I have not known a more intelligent, hard-working and dedicated group of people. I have learned a great deal from them in my years of service, and I thank them for this education.

Second, I owe deep thanks to the TUG office and our former Executive Director. I believe that he and the rest of this well qualified group of people devoted many cumulative decades of effort to move TEX forward in times when progress may have been much more difficult than it is today. I think that some of these people have neen treated in an unconscionable fashion, and I deeply regret that. These friendly and competent people were a considerable source of joy in my life, and I will miss them.

Third, I owe thanks to the TUG membership—including, of course, the non-dis-sonant members of the Board—for having had the opportunity to work with them over these past several years. Serving TUG has been an honor and a privilege. I have deeply valued the contributions to my understanding of TEX by many of these people, both TUG old-timers and some newtimers as well.

… and Some Not I owe no thanks, however, to the dissonant members of the Board. I find their actions ill considered, unfortunate and reprehensible, and will take at least some pleasure in the fact that I need not read their logorrheaic, self-satisfied diatribes any further. This will free hours and days of my life and—hopefully—rid me of the anger that their poor taste, narrow-mindedness and lack of consideration have cost me over the past several months.

Very Truly Yours David Ness 26 December 1990

To: Nelson Beebe

President, TEX Users Group

Date: January 16, 1991
Cc: Executive and Finance Committees

TUG Board of Directors Barbara Beeton Ron Whitney

Re: Resignation as TUG Vice President

Dear Nelson, I wish to submit my resignation as TUG Vice President, effective as soon as you appoint a replacement. As you know, I have been contemplating this action since the Board’s actions at Cork, but decided to delay acting until I felt TUG was in a stable enough position to avoid immediate collapse. I believe that I now have done all that I can to help achieve this goal and that it is now time for me to leave.

I agree with the points David Ness has made in his letter of resignation from the position of Treasurer. The Board’s actions at Cork concerning the Executive Director violated my sense of fairness and decency. The manner in which these actions were carried out has changed my view of the TUG governing body from one of a group working cooperatively together to build a strong organization to one dominated by factions and hidden agendas. I was disturbed by the total refusal on the part of the Board to consider the position of its Executive and Finance Committees and by the Board’s total rejection of the counsel of its past Presidents. I am tired of the amount of verbal and electronic mail sniping that has become a common response to all actions taken by the elected officials. I am distressed by the apparent abandonment of broad-based problem solving in the “new” TUG and its replacement with confrontation. I find the current climate of blame-casting and micromanagement destructive. As I cannot possibly take the additional energy away from my other duties that TUG now requires, I feel it best to step aside with the hope that someone else will be willing to take up the commitment.

In the opening paragraph of this letter, I said that I felt my responsibility was to remain in my position until the threat of immediate collapse had passed. I believe that this has now occurred, and that this is in large part due to the actions of the Executive and Finance Committees. The TUG office staff is talented, dedicated, and motivated. They are taking the steps necessary to preserve TUG and to forward its goals. They deserve and require the strong support of the Board. I would strongly urge the members of the Board to keep the stability of the organization uppermost in mind, and to avoid the always attractive temptation to change for change’s sake.

The Board has a major job upcoming in the summer meeting—that of restoring the sense of cooperation it trampled over in the Texas and Cork meetings. I urge the board to be moderate in its decisions. TUG retains a great deal of strength, which should be used as a basis for building and not discarded.

Like David, I would also like to request

TEX and TUG NEWS  Vol. 0, No. 0, May 1991  23

that this letter be printed in a forthcoming TUGboat. In part, I want to do so, to try and provide my perspective to the membership of TUG. One’s view of TUG varies dramatically, depending on what one’s connection is with the organization. I feel it important to separate in one’s mind the view of those interested in TUG for TEXrelated reasons and the view of those interested for organizational reasons. Although the organizational structure is rather unstable, I believe that the technical functions have been unimpeded—TUGboat and TeXHaX continue publication, the preparations for the annual meeting are ongoing, and the TUG office is functioning and continuing to provide day-to-day services to the membership.

I would like to crib another of David’s thoughts, and express my thanks to the present and former members of the Executive and Finance Committees, to the former Executive Director, and to the TUG office staff. I only wish that the rest of the Board and the membership as a whole could have witnessed the effort and attention that they have paid to TUG.

In closing, I would like to say that there is no doubt in my mind that every member of the board has the best interests of TEX and TUG in mind and believes that the actions they are taking are to further TEX and TUG. It is tragic that actions taken with such good intentions are resulting in so much upheaval and distrust.

February 13, 1991 Nelson Beebe President, TUG

Dear Nelson, This letter is to advise you of my resignation as Secretary of TUG. Since I have not discussed this decision with you, the choice of effective date is left to you. My preference is that the resignation be accepted immediately, but if you desire that I retain the office for a brief period, this is acceptable to me. Considering the events that TUG management has experienced since the Texas 1990 meeting, and the other resignations, I think it important to advise you of the reason for my decision. Of course, this correspondence may be distributed to the Board and to the membership. If distributed, I ask that the entire contents be used. I believe that the Board desires to change the composition of the Executive and Finance Committees. First, I am confident that the Board, as a body, feels that a change is essential to the future of TUG. I do not want to impede anything that will enhance the future of TUG. Over the years, I have always encouraged individual members to speak out and promote changes that they believed to be important to TUG’s growth and management. I still encourage constructive comment and action. Second, the Board’s approach to enhancing the future of TUG has been unacceptable to my standards of fairness, decency, and courtesy. This perspective is solely based on the manner the Board used to remove the past Executive Director. In my opinion, the person was removed for simply following the directions and guidance of the Executive and Finance Committees. Additionally, it appears to me that the Board was influenced by innuendo and was appealed to on the basis of personal prejudice. The embarassing realization is that I could have behaved as the majority of the Board; that is, the actions that have occurred lie within my concept of “ordinary” human behavior. I do not think that my not behaving in concert with the Board was extraordinary; it was more likely due to a difference in values and/or priorities. It is important that governing bodies have value and priority differences; when present in a community of respect, it is an effective way to provide representative government. As I do not feel there is a “community of respect,” my presence is an impediment rather than an aid. Finally, I want the best for TUG. And, I believe that all the Board members want the best for TUG. To my, my presence—considering my values and judgments—does not afford me the opportunity to contribute in my way toward the best for TUG. I am confident that Board activities will be more effective without me than with me.

The editor of TUGboat, Barbara Beeton, solicited commentary to David Ness’ letter of resignation. David’s comments are both elegant and are, I think to the point of some of TUG’s problems. The elegance is in his requirement to retain and assert his values. David and I do not agree on many things, but I know he respects my being and expression and I cherish the experiences we have shared. I hope each Board member and each TUG member will reflect on David’s comments, not in terms of just his position, but also his philosophy. In fact, the experience of working with the Executive Committee has been very rewarding to me and I think their efforts have been rewarding to TUG members.

I cannot end this correspondence without expressing my appreciation to the personnel at the TUG office, appreciation for tolerating and surviving this burst of growth by their employers. A fine group. And I hope the Board will be humane to Ron Whitney, a person who accepted a role to help TUG—it still confuses me why a person would enter this war zone.

Nelson, your interest in the future and potential of TUG has been pretty exciting to me. I am truly sorry that when the zeal and position connected, there was another problem that took priority. Yours truly,

Calvin Jackson

1991 TEX Users Group Election

After lengthy discussions over several years, the Board of Directors decided at a meeting held on March 2–3, 1991, that TUG would be best served by a Board whose members are elected by the full membership. To that end, an Election Procedures Committee was appointed to devise procedures which will result, this fall, in the election of an entire new Board. The committee has now completed its draft. At press time, the draft procedures await final approval by the Board.

The size of the Board will change; under the new rules there will be a maximum of 15 elected Directors. Several non-elected vice presidents—elected officers of major TEX user associations outside the U.S.—will also remain on the Board. Of the regular officers, only the President will be elected by the full membership; the other three—Vice President, Secretary, and Treasurer—will be chosen by the Board from among the elected Directors, to ensure that that these new officers are familiar with the amount of work expected of them. For this transitional period, the current Board will select the next President at their July meeting in Dedham.

Candidates for any open office may be nominated either by petition or from the floor at the annual business meeting. A petition must be signed by the candidate (to signify acceptance) and by two other TUG members. (A petition form follows.) The deadline for receipt of petitions for this year is September 1. A member nominated from the floor at the annual meeting must notify the TUG office in writing within seven days to accept the nomination.

A short biography of the candidate (not to exceed 300 words) and a candidate’s statement of purpose (about 100 words) should accompany the petition or acceptance letter. These statements will accompany the ballot sent to members, and may in some cases provide all the information that a member knows about a candidate.

Ballots will be mailed to all members about 30 days after the close of nominations. Marked ballots must be received no more than 5 weeks following the mailing; the exact date will be noted on the ballots.

Ballots will be counted by a disinterested party not part of the TUG organization. The results of the election should be available by the end of November.

All 15 regular positions on the Board will be up for election this yesr. This will make for a rather lengthy ballot. It is the Board’s hope that all members of TUG will participate actively, either as candidates or by supporting the candidacy of others, and certainly as thoughtful voters.

Barbara Beeton

for the Election Procedures Committee

Nomination for 1991 TUG Board of Directors Election

Only current (1991) TUG members are eligible to have their names appear on this petition. The signatures of two (2) members are required in addition to that of the nominee. Type or print names clearly, exactly as they appear in the most recent TUG membership list or on a TUG mailing label; new members should enter the name which they used on their membership application form. Names that do not exactly match the TUG records will not be accepted as valid.

The undersigned TUG members propose the nomination of:

Name of nominee  (signature)  (date)

for the position of Member of the TUG Board of Directors, for a term beginning January 1, 1992.

Members supporting this nomination

Nominated by  Signature  Date

(please print)

Return this petition to the TUG office (FAXed submissions will also be accepted). Petitions must be received in the TUG office no later than September 1, 1991.

TEX Users Group FAX: 401-751-1071 Nominations for 1991 Election P.O. Box 9506 Providence, RI 02940-9506 U.S.A.

TEXhax Tidbits

TeXhax Digest       Sunday,  March 17, 1991      Volume 91: Issue 013
Moderators: Tiina Modisett and Pierre MacKay
%%% The TeXhax digest is brought to you as a service of the TeX Users Group %%%
%%%       in cooperation with the UnixTeX distribution service at the       %%%
%%%                      University of Washington                           %%%
Today's Topics:
                help for importing postscript files in latex
                          Wanted: Armenian in TeX
                                 font names
                                    detex
                   TUG Conference Proceedings Summaries
                    AT&T patent => X will not be free.
            Standardization of TeX names for Adobe PostScript fonts.
          Re: Standardization of TeX names for Adobe PostScript fonts.
-----------------------------------------------------------------------------

Look familiar? To the thousands of TEX users who have network mail access, reading TEXhax is a part of their daily or weekly (well, ok, monthly) ritual. TEXhax is an electronic newsletter moderated by Tiina Modisett and Pierre Mackay at the University of Washington/Seattle. However, many TUG members do not have access to TEXhax; and many TEXhax users are not TUG members. And many users of TEXhax are not aware of the fact that TUG is in large part responsible for the continued availability of TEXhax: TUG contributes some $10,000 per annum to maintaining this vital open news forum.

A regular column in this newsletter, summarising the main items of interest in recent issues, would allow the entire TUG membership to benefit from the tremendous amount of information which flows electronically via TEXhax. It would bring back to TUG members some of their investment in supporting TUG, and would also bring a better appreciation of TUG’s involvement in providing services to the TEX-using community, regardless of their membership in the user group.

A volunteer or volunteers for this work would certainly be welcome. Send an outline of what you would do, what you would like to see in this space, to tug@math.ams.com, and mark the Subject line TeXhax column. Suggestions for names for the column are also invited.

Past volumes of TEXhax are available on floppy diskette in Jon Radel’s collection of TEX- and -ware for Microcomputers. This collection is now being distributed by the TUG office in Providence. Contact the TUG office at tug@math.ams.com for a full catalogue or further details.

Update: 12th TUG Annual Meeting July 15–18, 1991

The 1991 Annual Meeting of the TEX Users Group is shaping up well, and the promotional brochure should have reached you all by now. Here are some updates on the program:

1. There will be a Hospitality Suite this year (an oft-made suggestion in the past). Look for the signs at the Dedham Hilton when you arrive on Saturday or Sunday, July 13–14, 1991 (8 a.m. to 10 p.m.).

2. Alan Hoenig will present a 30-minute Introductory Talk to TEX, which will be of particular interest to newcomers to TEX to TUG. The talk will take place Monday, July 15th, 8:30 a.m. in the main ballroom.

3. There will be a Keynote Speaker (to be announced) on Monday, July 15th, 9:15–10:00.

4. There will be two Networking Lunches, one on Monday (as announced) and a second one for Thursday. Return your form to the TUG office as soon as possible. Sign up early and be counted!

5. There will be three workshops on each of the first three days. Sign up at the registration desk. Correction: The workshop “Modifying manmac to Suit the Publisher” will be led by Dan Olson, not Mimi Lafrenz; both are from Electronic Technical Publishing, not the firm indicated.

6. On Wednesday, July 17th, have Lunch with the Board. Members of the board of directors will be making themselves available so you can talk to them. Talk about TUG as an organisation, make suggestions for the various committees to consider, volunteer your time and/or expertise, find out what it takes to be a board member.

7. On Thursday, July 18th, to close off the four days, we will have an Historical Round Table. Come meet some of the “names” in the TEX community; listen to reminiscences anecdotes and stories about the development of TEX, early early implementations on various systems, and how some of us became involved with TEX — not everyone was from the Math or Systems Engineering departments! Some come from English, linguistics, office management, and so on.

8. A LISTSERV address has been established at Texas A&M (thanks to Tom Reid) for posting queries to be used in the annual Questions and Answers session at the meeting. Send queries to: tug-q@tamvm1.tamu.edu or tug-q@tamvm1.bitnet. Remember: you won’t get an answer to your posting — this is just a place to send queries for use at the July meeting.

9. There will be a PC and a Mac set up for a software exchange, with lots of software available for copying. If you have macros you’d like to share, bring them and add them to the collection. Bring your own disks.

10. And here’s another way to share your ideas — I’m calling it Admin.TeX. Bring samples (hardcopy of input/output) of office-related documents, such as letterheads, memos, forms (FAXes, invoices, whatever), anything that you use in an administrative/office context, produced with TEX or L^{A}TEX. We’ll post it, and you can show off what you do with TEX. Share your expertise with colleagues; find something that you’ve always needed. Try to keep the input/output to a page or two; bring the macros too if you want, and put them on the software exchange machines.

11. New this year — an on-site TEXpert! A variety of people will be donating their time to sit at a machine, and try to help you solve your problems. Sometimes all it takes is a fresh set of eyes to look at your code; remember though that there’s only so much you can ask of someone who’s got a lot of questions lined up …

12. Program Changes: “The Realities of TEX as a College Textbook Publishing Program” (Mona Zeftel) has been replaced by “The Five C’s: A Guide to Successful Publication Using TEX” (Colleen Brosnan, Prentice-Hall); “TEX’s Role in a Computer Integrated Administration (CIA) Environment” (H. Petersen) has been withdrawn.

So! Make sure you send in your registration form now for this year’s Annual Meeting! Pass the brochures around, make copies of the form, and come and join us for a truly dynamic meeting. Don’t forget: everything takes place at the Dedham Hilton, July 15–18, 1991. Plan to arrive on the weekend and take in some of the sights in and around Boston; pop into the Hospitality Suite and see who’s there; take a swim in the pool or a walk around the park-like grounds of the hotel. Make this an enjoyable conference — relax and talk TEX!

Christina Thiele

TUG ’91 Program Coordinator May 1991

“Hey — it works!”

Are there little tricks you use that you don’t even think about? Tricks that you figure are too trivial, or too inelegant to be of interest to other TEX users? A lot of TEX and L^{A}TEX users are just looking for a quick fix now and then; and a lot of us who’ve been using the programs for a while have our favourite little tricks.

Send in some of your best samples of sleight of hand … or code. Consider this a forum for sharing and exchanging helpful hints — but they’d better work! This isn’t a test site for your almost-ok ideas — test the thing and make sure it’s bug-free. And one more thing: items show in this column do not pretend to be the best or all-purpose solution to little problems. They show one way to deal with a specific problem in a specific context. Don’t attack a hack — send in a better one!

Here are a few of my own offerings, which may be howlers to some, but “Hey — they work!”.

Some applications for \llap

The following would work with either plain TEX or L^{A}TEX; I use macro packages which are refinements and variations on the basic manmac file.

Blank Boxes

Some might view this as a throw-back to the typewriter era … but the TEX control sequences \llap (and \rlap) can come in awfully handy (see the TEXbook, pp. 82–83; see also pp. 189, 340–341 for some more suggestions). And if you add \raise and \lower, then there are quite a few things you can do, along with \kern to do the fine tuning.

Essentially, these are commands to “back up” or “go forward”, to “go up” or “down”, and then move a shade to the left or right.

One item I once needed was a small box (in a semiotics paper, for some reason). L^{A}TEX users have $\Box$, which yields □ ; but plain TEX users have to build them from scratch. One way to get a box is to lap a math mode \sqcup back over a math mode \sqcap, like this:

   \def\blankbox{$\sqcup$\llap{$\sqcap$}}

and you get this: ⊔ ⊓ . Note that it’s a bit different from the L^{A}TEX box.

Thorns: upper and lower case, roman and italic

My background is linguistics, phonetics in particular. Sometimes when you need a character only once or twice, a quick solution is called for. Here’s one for a lowercase “thorn” (thorn appears in Scandinavian texts and in Old English, for example). For the lowercase thorn, print a p on top of a b, using \llap and you get: b p (this also works if you’re in an italic group: b p . Yes, there’s that little serif from the p, but if your alternative is to hand-draw the character, consider using this rough replacement:

   \def\thorn{b\llap{p}}

If you need a capital thorn, it gets a bit strange: define a very large font to get a big “el” to serve as the “stem” of the letter l; then back a capital O onto it:

   \font\bigl=cmr10 scaled\magstep2%    %% for a large roman stem
   %% UPPER CASE Thorn:
   \def\Thorn{{\lower.47ex\hbox{\bigl l}}\kern.46em{\llap{O}}}

So here’s a capital Thorn (roman style): lOis is it.

And some more …

h\llap{\lower2ex\hbox{\u{}} is for an h with a breve sign underneath (I think it was for some kind of a voiceless velar fricative), which an author needed: h̆ .

Another paper needed some very, very small equals signs between characters. Using cmr5, the equals are still too large = and would need to be raised. So, looking for something that’s even shorter, you come to the hyphen. Backing up with \llap isn’t enough, because that would just overprint, so you have to raise one of the hyphens:

   \def\qq{-\llap{\raise.3ex\hbox{-}}}

and--then--use--it … well, it’s got a very restricted usage.

But the point is, you can create all sorts of make-do characters by combining existing ones and controlling their position, relative to one another. And in a pinch, they’ll do the job.

⊔ ⊓  ⊔ ⊓  ⊔ ⊓

Perhaps this playing around with some fairly simple commands will be an incentive to the hundreds and thousands of TEX and L^{A}TEX users out there. Send in your own favourites. Explain the context (e.g., mainly in text, mainly in math, something specific to a particular L^{A}TEX environment). Let’s make this an exchange forum for all users of all programs.

See the inside front cover for information on submissions.

Upcoming Events

June 13 Stockage de documents (last in a lecture series “Les documents électroniques”) Paris, France Contact: Mme Tourteau, EDF–GDF/STI, 21, rue Joseph Bara, F–92132 Issy-les-moulineaux +33 1 44 26 87 32

July 15–18 TUG ’91: “TEX Inroads into Publishing” Dedham, Mass. (suburban Boston)

TUG Office 401-751-7760 Fax: 401-751-1071

tug@math.ams.com

Sept.23–25 EuroTEX 91: 6th European TEX Conference Paris, France Contact GUTenberg, B.P. 21, F–78350 Jouy en josas +33 1 34 65 22 32 Fax: +33 1 34 65 20 51

gut@irisa.irisa.fr

Sept. 26 GUTenberg ’91 Conference: “Technical and Scientific Publishing” Paris, France Contact: GUTenberg (see address above)

{Sept. 30– \atop Oct. 3} Computer Publishing Conference San Jose, California Contact: Seybold Publications 213-457-5850

Oct. 1–3 Desktop Publishing in Astronomy and Space Sciences Strasbourg, France Contact: André Heck, Observatoire Astronomique, 11, rue de l’Université, F–67000 Strasbourg +33 88 35 82 22 Fax: +33 88 25 01 60

heck@frccsc21.bitnet

Oct. 15–16 RIDT 91: The 2nd International Workshop on Raster Imaging and Digital Typography Boston, Mass.

Contact: Robert A. Morris 617-287-6466

ridt91-request@cs.umb.edu

Note: Course information will appear in TUGboat 12, no. 2. For more information on conferences or courses, contact the TUG office (address on inside front cover).

TEX and TUG NEWS Table of Contents

Editorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 jeep: A General Purpose Style File

Peter Stewart . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 News from Around . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

TEX in the Soviet Union

Theo Jurriens . . . . . . . . . . . . . . . . . . . . 5

New LISTSERV for Greek TEX Users

Yannis Haralambous . . . . . . . . . . . . . . . . 6

The Truth about TEX

Christina Thiele . . . . . . . . . . . . . . . . . . 6

A Message from the Board of Directors . . . . . . . . . . . . . . . . . 7 New Focus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Some Changes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

Summary of TUG Board of Directors Meetings . . . . . . . . . . . . . 11 September 1990 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 March 1991 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

Communications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 1991 TEX Users Group Election . . . . . . . . . . . . . . . . . . . . . . 25 Nomination for 1991 TUG Board of Directors Election . . . . . . . . . 26 TEXhax Tidbits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 Update: 12th TUG Annual Meeting . . . . . . . . . . . . . . . . . . . 28 “Hey — it works!” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

Some applications for \llap

Christina Thiele . . . . . . . . . . . . . . . . . . . . . . . 30

Upcoming Events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

Eight Great Reasons to Join the TEX Users Group

1. TUGboat All members receive a subscription to TUGboat, The Communications of the TEX Users Group. Published four times a year and averaging 150 pages per issue, TUGboat contains solutions to typesetting problems, bug fixes, and information on new developments in TEX, L^{A}TEX, AMS-TEX and other macro packages. Each issue has a calendar of coming events and various resource lists to help readers find information about existing macros, drivers, and TEX publications.

2. Books and Software A “one-stop shop,” TUG carries most TEX macro packages as well as TEX publications and products. The Radel collection of public domain TEX software for microcomputers is now available through TUG, too.

3. Membership Networks TUG maintains and distributes a complete membership list annually. Members are listed alphabetically, geographically and by institution, as well as by the computers and printers they use. These lists facilitate easy member-to-member communication within TUG’s worldwide network.

4. Information Referrals TUG can connect you with site coordinators — specialists in implementations of TEX on various computer architectures — as well as other resource volunteers. These experts are available to answer your questions about TEX, TEX-related software and other technical matters.

5. Annual Meetings TUG’s Annual Meetings bring TEX users together to learn the latest in TEX applications and innovations through seminars, talks and informal gatherings. They offer an opportunity to connect with other TEX users at every level. Each conference is informative, but always informal and friendly — the perfect forum for valuable interaction and exchange of ideas on TEX.

6. TEX Worldwide When you belong to TUG you connect with other TEX user associations around the world. Information about their activities is published regularly in TUGboat and reciprocal membership arrangements are available with some of the European groups.

7. TEX Training Each year TUG offers seminars and classes for beginner, intermediate and advanced TEX users throughout North America and Europe. TUG will custom design and conduct on-site courses to meet your company’s specific needs.

8. Discounts

▷ Discounted Annual Meeting fees ▷ Student rates for membership and courses ▷ Discounts on the purchase of selected TUG publications ▷ NEW! Discounts for all members on TUG-sponsored courses in TEX and TEX-related subjects